ykxmileyx:

A student simplifies (6b + 4r) – (2b + r) and says that the result is 4b + 5r. Explain the error and describe the correct steps to simplify the expression.

jimthompson5910:

Let's say two students have only books and rulers in their backpacks If one student has 6 books in their backpack, we can shorten that to saying 6b b = 1 book 6b = 6 books If this same student had 4 rulers, then we write 4r r = 1 ruler 4r = 4 rulers This leads to the expression 6b+4r indicating 6 books and 4 rulers The expression (2b+r) means the other student would have 2 books and 1 ruler (6b+4r) - (2b+r) is what happens when we subtract the two backpacks Specifically we subtract the number of books together 6 books - 2 books = _____ books and do the same for the rulers too 2 rulers - 1 ruler = _____ rulers Or in shorthand we write 6b - 2b = _____ 2r - 1r = _____

jimthompson5910:

Hopefully that all makes sense

supie:

Yeah, that makes sense jimthompson5910 What i did was: \[6b + 4r - (2b + r)\]\[ 6b + 4r - 2b - r\]\[(6b - 26) + (4r - r)\]\[ 4b + 3r \] sooo idk

jimthompson5910:

For some reason I wrote the wrong expression for the rulers portion I should have said 4 rulers - 1 ruler = _____ rulers which shortens to 4r - r = ____

jimthompson5910:

you have the correct steps @supie though you wrote 6b-26 instead of 6b-2b

supie:

Oh ok thanks bro.

jimthompson5910:

no problem